write an equation of the line that passes through point (4,7) and is perpendicular to the line whose equation is 2x+3y=4

Answer by stanbon(75887)

write an equation of the line that passes through point (4,7) and is perpendicular to the line whose equation is 2x+3y=4
---
slope-intercept form of the given line: y = (-2/3)x +(4/3)
Notice the slope is -2/3
----
Any perpendicular line must have slope = 3/2
------
Find equation of line with m = 3/2 passing thru (4,7)
Form: y = mx + b
7 = (3/2)4 + b
7 = 6 + b
b = 1
----
Equation:
y = (3/2)x + 1
=====================
Cheers,
Stan H.
=====================